A particle of mass M1 moving with initial velocity  is incident on a stationary particle of mass M2.After collision, M1 was deflected through an angle  and M 2 an angle .If the velocity of the particles after collision were  and  respectively.Show that for an elastic collision;

(a)    = 2


(b)   =2 +



(c)    =


1
Expert's answer
2021-02-23T04:04:53-0500

Given,

Masses of the particles are "M_1" and "M_2" and their respective velocities after the collision are "v_1" and "v_2" and initial velocity of the mass "M_1" is "v_o"

Let after the collision the deflection angle be "\theta" for both,

Now, applying the conservation of linear momentum,

"\Rightarrow M_1v_0=M_1v_1\cos{\theta}+M_2v_2\cos{\theta}"

"\Rightarrow M_1v_1\sin{\theta}=M_2v_2\sin{\theta}"

Now, substituting the values,

"\Rightarrow M_1v_0=2M_2v_2\cos{\theta}"

Now, again,

"v_2=\frac{M_1v_0}{2M_2\cos{\theta}}"